• Security incident: ISF was recently accessed by intruders. Please change your password, and change it anywhere else you used it. Read more

VFF to Appear on Rational Alchemy

Ahh, Ok, I think I see what I meant, and I misspoke. Statistics are not a very strong suit for me (we always have a statistician on hand for the challenges.) I apologize. You were right to ask about those, because they make no sense.

I believe Jeff Corey's stats are correct.
 
But Anita's p value was .23, nowhere near that. More like 1 in 4.
Wrong. Her P-Value was .0567, as explained below:

I calculated the probabilities for all of the results that would be
inferior to Anita's result, which was getting two persons correct and
one location correct, and then subtracted the sum of those
probabilities from 1. So, I calculated the probabilities of (a)
getting none of the persons correct; (b) getting only one person
correct, with the location either incorrect or correct; or (c) getting
two persons correct, but with both locations incorrect.

For (a), it's a simple matter of multiplying 5/6 * 5/6 * 5/6, which
equals 125/216. The reason for this is that the probability of getting
each person incorrect in each of the three trials is 5 out of 6. If
that occurs, no calculation is necessary for location.

For (b), it's a little trickier because the correct person could be
chosen in any of the three rounds, but we still don't need to do a
separate calculation for location because, whether the location is
correct or incorrect, the result (one person correct, but with the
incorrect location or one person and one location correct) would be
inferior to Anita's result. Thus, the overall probability of getting
one person correct is 3 * 1/6 * 5/6 * 5/6, which equals 75/216.

For (c), it's trickier still because now the probability of getting
both locations incorrect after getting two of the three persons
correct must be accounted for. However, that's simply a matter of
factoring in the 1 in 4 probability of that happening. Again, there
are three ways of getting two persons correct (correct in Rounds 1 and
2, correct in Rounds 1 and 3, or correct in Rounds 2 and 3). Thus, the
probability getting two persons correct, but both locations incorrect,
is 3 * 1/6 * 1/6 * 5/6 * 1/4, which equals 15/864.

Summing 125/216 + 75/216 + 15/864 = 815/864. Subtracting that sum from
1 = 49/864, or 5.67%.
 
If a dowser claims a 100% success rate, is tested and can't meet the protocol of the challenge, they have failed.

Should they be retested? Resources permitting, of course they should.

Why? Because belief in dowsing still persists. All we can do as skeptics is say "Show us, and we'll believe." If we stop saying "Show us," we lose.
100% agreed, though note...you've not attributed this retesting to "pretty close" or other ambiguous reasons as you have with VisionFromFeeling.

To be clear: I'm not faulting the IIG or their conclusions. Their stated results are exactly correct. There is NO ambiguity in the results at all, as is proper. However, what I observed was interesting compared against other challenges. And if VFF feels she'd like to be tested again, I think that's fine.
There was nothing ambiguous.
Then please stop introducing ambiguity, with "pretty close" or partial credit as you and Nigel have done.

From the IIG Site:
Anita Ikonen Fails the Preliminary Demonstration said:
Anita did not succeed. She chose incorrectly on trial one, chose correctly on trial two, and chose incorrectly on trial three.
There is no "pretty close", no partial credit for trial three.
 
Okay, I just finished listening to the RA podcast and read through most of these threads. I'm sick of the odds! She failed the test. Done!

Some of the things I am interested with the RA podcast are...

Anita said she does not "see" the kidney but feels its presence. She said that when the person has both kidneys they feel different than someone who has an imbalance with only one kidney. This was not clear to be before, so now I understand why she felt uncomfortable having to choose a side. The IIG insisted on her choosing the side otherwise it would have been too easy to just pick out a person.

I do think it was very interesting that she choose the left side each time as she admitted that she knew that is the side that is more often missing.

But she did agree to the protocol. I didn't really like that the hosts on RA were trying to give her the third trial just because she picked the right person. One in six were the odds of getting the correct person if you just look at the single trial. (yes I know I'm talking about odds here) and that does not seem like much of anything, yet the hosts made it sound like a really big deal.

I had heard that the podcast slammed the IIG. I didn't hear that, they did seem to be downplaying all the work that it took to put this together. Jeff says that it should not be a problem getting re-tested, he mentions that she could try for the JREF challenge and also the Tampa Skeptic challenge. Then he says that maybe the IIG would re-do the challenge without the money award, just for practice (something like that). What?

Finding people missing kidneys and arranging everything they did was very time consuming and difficult. No way was that something they threw together and Jeff made it sound like it would be something simple to do again. If Jeff was thinking about a easy test then he should have said so, and IF an easy test exists that Anita would agree to then she should set it up herself at home and complete it.

Why would the IIG be bothered with her again? She failed (yes she was nice about it) but she still failed.

I also listened to Jeff and Anita's interview that exists on her site.

Jeff says he does not believe that she has powers, he is clear about this and says it several times (both on the interview and the podcast). Yet he also is clear that just because she failed does not mean she does not have that ability. He says that maybe at that moment her skills failed her, that throwing reindeers off the roof and they don't fly does not mean they can't fly. They might not want to fly at that moment, or maybe that specific reindeer can't fly. WHAT?

This sounds so much like Gary Schwartz and his white crow comment. (just because you have never seen a white crow does not mean they don't exist, he likens this to finding some evidence of a person who can communicate with the dead).

If you follow Jeff's logic then it is impossible to ever prove someone can not perform whatever claim they fail. If they fail the test, then just say "their powers disappeared for that moment and there is no control over the powers". It reminds me of people who would argue endlessly that just because we see the sun rise in the morning it might not rise again tomorrow morning. (yes, I do understand it does not rise, the Earth moves and it appears to rise, calm down people)

At some point you have to STOP and say. What is your claim, with what accuracy? How can we test it? And that's it! When they fail it is silly to say, well maybe something happened, you didn't feel strong at that moment. No excuses! What is the claim, with what accuracy and test it. DONE.
 
I was also surprised to hear that Anita talks to dead people and they are the ones starting the conversation. Hello!

Where do you get "I can sense missing kidneys" to "dead people talk to me while I watch TV"?

She wants Jeff to take her ghost hunting. Ahhhh No!

Lets remember Kitty telling us about unicorn wings. When you discuss whether or not unicorns have wings you fall into the woo trap of saying there are unicorns. FIRST you need to prove that unicorns exist before you start talking about their wings.

Before you can test whether or not Anita can sense ghosts, you need to prove there are ghosts.

She wants to go to places where there have been reported cases of ghosts sited. Where on Earth has there NOT been a ghost sited? OR better yet, where can we say for certain that it is NOT possible for a ghost to be hanging out at this moment, or yesterday or last month or ten seconds from now or....? How can you pick a place when the woo can just say, "yeah well, maybe he didn't die here but he used to visit this place as a kid and that's why I see his ghost here" OR "as a child the ghost used to read about this place in a book and that's why his ghost is present here". It's endless.

Proving ghosts is like proving prayer heals. There is no way to set up a test with any kind of control. Woos can spend the day saying they see dead people over your right shoulder, so can I. What does that prove, nothing.
 
100% agreed, though note...you've not attributed this retesting to "pretty close" or other ambiguous reasons as you have with VisionFromFeeling.


Then please stop introducing ambiguity, with "pretty close" or partial credit as you and Nigel have done.

From the IIG Site:
There is no "pretty close", no partial credit for trial three.

I don't think I've ever wavered from the fact that she completely failed.
 
I was also surprised to hear that Anita talks to dead people and they are the ones starting the conversation. Hello!

Where do you get "I can sense missing kidneys" to "dead people talk to me while I watch TV"?

She wants Jeff to take her ghost hunting. Ahhhh No!

Lets remember Kitty telling us about unicorn wings. When you discuss whether or not unicorns have wings you fall into the woo trap of saying there are unicorns. FIRST you need to prove that unicorns exist before you start talking about their wings.

Before you can test whether or not Anita can sense ghosts, you need to prove there are ghosts.

She wants to go to places where there have been reported cases of ghosts sited. Where on Earth has there NOT been a ghost sited? OR better yet, where can we say for certain that it is NOT possible for a ghost to be hanging out at this moment, or yesterday or last month or ten seconds from now or....? How can you pick a place when the woo can just say, "yeah well, maybe he didn't die here but he used to visit this place as a kid and that's why I see his ghost here" OR "as a child the ghost used to read about this place in a book and that's why his ghost is present here". It's endless.

Proving ghosts is like proving prayer heals. There is no way to set up a test with any kind of control. Woos can spend the day saying they see dead people over your right shoulder, so can I. What does that prove, nothing.

Unless... I can see them too. And while I've slept in a few haunted houses, and even in a dead man's bed and have never seen anything evenly remotely "ghostly," maybe I will some day.

Of course, then I have a problem. How can I tell a ghost from an alien?
 
If Anita had to pass the test 100% then she wouldn't be the focal point,and neither would JREF,so why all the bother with her? Are the members aware of this obvious fact?
 
Are the directors of JREF and the members addicted to Anita? Please seek some professional medical help if need be the case.:D

I think you all are a bit wacky ;)
 
Well, if you're with Anita...the aliens are the ones who aren't notable figures in American history. :D

For those who are new to the VfF game. desertgal is alluding to the fact that VfF believes that she, herself is an alien. So, even if he went ghost hunting with her and they didn't find a ghost, he would still be with an alien.

Ward
 
Im willing to bet she'd find ghosts. SHe could come to my house and talk to my monkey ghost. Lisa has a demon monkey also.

I just feel its kinda like spoon bending. There are easier ways to tell if someone has a kidney. Now get a FUN COOL supernatural power, like invisibility or the ability to climb walls like a spider.

I want to test someone like that (Maybe Mac King could try for the invisibility)
 
For those who are new to the VfF game. desertgal is alluding to the fact that VfF believes that she, herself is an alien. So, even if he went ghost hunting with her and they didn't find a ghost, he would still be with an alien.

Your messing with me right? She doesn't really.
 
Your messing with me right? She doesn't really.

I am sorry to say that you are not being messed with. She believes she is an incarnation of a dwarf star somewhere near Arcturus, presently in human form, and somewhere in the neighbourhood of 350 years of age, give or take a few decades. This is the supposed source of all her "magical" abilities. It is also this alien genetics that allows her to live on nothing but air for a week or two with no ill effects. In fact, it supposedly makes her feel better than eating does. I could go on, and on, and on, but the StopVfF website does a better job of succinctly listing all the claims than I could.
 
Unless... I can see them too. And while I've slept in a few haunted houses, and even in a dead man's bed and have never seen anything evenly remotely "ghostly," maybe I will some day.

Of course, then I have a problem. How can I tell a ghost from an alien?

Careful, Jeff. Senex has already volunteered to do testing of ghost sightings at haunted hotels, and if you make the same offer, he might get jealous. :D

The ghost will be the semi-transparent apparition floating nearby. The alien will be the plus-sized blonde pointing out the ghost. :D
 

ISF - Join now!

Every member here is approved by hand. No bots, no spam, just people who care about evidence and honest debate.

Membership is free!

Create your free account

Back
Top Bottom